# Density Estimation via Measure Transport: Outlook for Applications in the Biological Sciences

### v2

@@ 3 Basics of Measure Transport
Recall that we are concerned with the problem of estimating an unknown probability density function characterizing some given data set of samples or observations, for instance, gene expression data. Denote the corresponding unknown probability measure by $\nu_{\pi}$ and call it the target measure. On the other hand, let $\nu_{\rho}$ be a probability measure amenable to computations, for example, the standard normal $\mathcal{N}(0,I)$, so one can easily draw samples from it and evaluate its density $\rho$. Call this the reference measure. Now, let $\mathcal{{X}}=\left(\mathcal{{X}}_{1},\mathcal{{X}}_{2},\dots,\mathcal{{X}}_{m}\right)$ and $\mathcal{{Y}}=\left(\mathcal{{Y}}_{1},\mathcal{{Y}}_{2},\dots,\mathcal{{Y}}_{m}\right)$ be random vectors with probability distribution functions $F_{\mathcal{{X}}}$ and $F_{\mathcal{{Y}}}$, respectively, such that $\mathrm{d}{{F_{\mathcal{{X}}}}}=\nu_{\rho}$ and $\mathrm{d}{{F_{\mathcal{{Y}}}}}=\nu_{\pi}$. The random vectors $\mathcal{{X}}$, $\mathcal{{Y}}$ assume, respectively, values $x,y\in\mathbb{R}^{m}$ (also denoted by $x\sim\nu_{\rho}$ and $y\sim\nu_{\pi}$). Then if $X=\{x\in\mathbb{R}^{m}\,:\,x\sim\nu_{\rho}\}$ and $Y=\{y\in\mathbb{R}^{m}\,:\,y\sim\nu_{\pi}\}$ are sets of samples formed from the values of the random vectors $\mathcal{{X}}$ and $\mathcal{{Y}}$ (i.e., the elements in $X$ are distributed according to $\nu_{\rho}$, and those in $Y$ are distributed according to $\nu_{\pi}$) the transport map should satisfy $\nu_{\pi}(Y)\ =\ \nu_{\rho}(T^{-1}(Y))$.
@@
Thus, given the sets of samples $X$ and $Y$, our goal is to learn (or compute) the transport map $T$ and use it later as a predictor for the values of $\mathcal{{Y}}$ (and therefore an estimator of $\nu_{\pi}$). The set $Y$ consists of samples (usually taken from experiment) of the random vector $\mathcal{{Y}}$ distributed according to an a priori unknown target probability measure $\nu_{\pi}$, while the reference probability measure $\nu_{\rho}$ is a priori known (e.g., Gaussian) and the set $X$ is generated from samples of a random vector $\mathcal{{X}}$ distributed according to $\nu_{\rho}$. Under the assumption that one can compute such a map $T$ transporting between the target and reference measures, the framework outlined herein and depicted in Figure 2 yields a method for approximating the desired unknown density function that characterizes the given sample data set. Then, the central problem is to compute, or learn, such a map $T$. This problem is further addressed in Section 4.
@@
Upon learning such a map $T$ from data, evaluation of the (unknown) density $\pi$ at a given (new) sample $y\sim\nu_{\pi}$ can be done via equation (4) as, by construction, all the functions on the right-hand side of equation (4) are known and can be evaluated. Moreover, because the reference measure is chosen to make sampling from it an easy and efficient task, sampling from the target measure can be done efficiently by drawing samples from the reference and transporting them to the target via the learned map $T$. Thus, probabilistic modeling, inference, and statistical analyses in the target space are enabled by such measure transport framework.
@@ 4 Transport Map Computation
To compute transport maps, we use software [28, 3] made available by the MIT Uncertainty Quantification (MUQ) Group [27]. The reader is referred to the original publications [10, 24, 40, 2] for complete descriptions of the numerical methods and software libraries. A summary is provided next, to make the present manuscript self-contained.
Let $\nu_{\rho}$ and $\nu_{\pi}$ be two (probability) measures defined on the algebra of Borel sets of $\mathbb{R}^{m}$, $m\geq 1$. Suppose that $T:\mathbb{R}^{m}\rightarrow\mathbb{R}^{m}$ is a transport map pushing forward $\nu_{\rho}$ to $\nu_{\pi}$. That is, it satisfies the equation
$\nu_{\pi}$ $=$ $T_{\sharp}\nu_{\rho}$ (5)
for $T$. By definition (see Section 3),
$\nu_{\rho}$ $=$ $(T^{-1})_{\sharp}\nu_{\pi}$ (6)
@@
and one has that $y_{i}=T_{i}(x_{1},x_{2},\ldots,x_{m})$, where $y=(y_{1},y_{2},\ldots,y_{m})\sim\nu_{\pi}$ and $x=(x_{1},x_{2},\ldots,x_{m})\sim\nu_{\rho}$. Conversely, the inverse transport map $T^{-1}$ of $T$ has $m$ components
@@
and one has that $x_{i}=(T^{-1})_{i}(y_{1},y_{2},\ldots,y_{m})$.
A transport map $T$ is computed via the solution of equation (5) and a minimization problem with cost or objective function $C_{\rho}(T)$ generally defined as
$C_{\rho}(T)$ $=$ $\int_{\mathbb{R}^{m}}c(x,T(x))\,\nu_{\rho}(\mathrm{d}{{x}})\,.$ (9)
The structure of the map $T$ depends on $\rho$, $\pi$, and $c(\cdot,\cdot)$. Many types of maps (not necessarily optimal), or couplings, exist between measures [44, 36]. In this study we consider the use of triangular transport maps, as they enable practical numerical computation of a transport map and its inverse [24, 2].
@@ 4.1 Triangular Transport Maps
A transport map $T:\mathbb{R}^{m}\rightarrow\mathbb{R}^{m}$, $\nu_{\pi}=T_{\sharp}\nu_{\rho}$, is triangular if
$y_{i}$ $=$ $T_{i}(x_{1},x_{2},\ldots,x_{i}),\quad i=1,\ldots,m.$ (10)
In other words, the $i$-th component $T_{i}$ of a triangular transport map $T$ only depends on the values $\{x_{j}\}_{j=1}^{i}$ of the first $i$ random variables $\{\mathcal{{X}}_{j}\}_{j=1}^{i}$. Thus, the map components are
$T(x)$ $=$ $\left[\begin{array}[]{l}T_{1}(x_{1})\\ T_{2}(x_{1},x_{2})\\ T_{3}(x_{1},x_{2},x_{3})\\ \qquad\vdots\\ T_{m}(x_{1},x_{2},\ldots,x_{m})\end{array}\right]\,.$ (16)
Conversely,
$x_{i}$ $=$ $(T^{-1})_{i}(y_{1},y_{2},\ldots,y_{i}),\quad i=1,\ldots,m.$ (17)
To ensure positivity of the densities (3) and (4), each map component $T_{i}$ in equation (10) must be differentiable with respect to $x_{i}$, and it must hold that $\partial_{x_{i}}T_{i}>0$, for every $i=1,\ldots,m$. We refer to the latter as the class of increasing triangular maps. Similarly, each component $(T^{-1})_{i}$ (17) of the inverse map $T^{-1}$ of $T$ must satisfy said properties with respect to the variables $y_{i}$ [4, 24]. Furthermore, when both $\nu_{\rho}$ and $\nu_{\pi}$ are absolutely continuous with respect to the Lebesgue measure $\mathrm{d}{{x}}$ on $\mathbb{R}^{m}$ with densities $\rho$ and $\pi$, respectively, if $c(x,T(x))=||x-T(x)||^{2}$, where |||||$\cdot$||| | ⋅ | | is the Euclidean distance on $\mathbb{R}^{m}$, and $\nu_{\rho}$ is Gaussian, instead of (9) one may seek to minimize the relative entropy, or Kullback-Leibler divergence,
$\mathrm{D_{KL}}(\nu_{\pi}\,||\,\nu_{\rho})$ $=$ $\int_{\mathbb{R}^{m}}\pi\log(\pi/\rho))\,\mathrm{d}{{x}}\ =\ \mathbb{E}_{\pi}(\log(\pi/\rho))\,,$ (18)
as $C_{\rho}(T)\leq 2A\,\mathrm{D_{KL}}(T_{\sharp}\nu_{\rho}\,||\,\nu_{\rho})$, where $A>0$ is a constant [31, 19, 4]. Under the assumption that the map $T$ is differentiable or even smooth (cf. formula (41) in Appendix A), one may also look at the problem of computing an optimal transport map $T$ as a solution of the optimization problem $\arg\min_{T}C_{\rho}(T)$, or indeed
@@
(conversely, $\arg\min_{T}\mathrm{D_{KL}}(\nu_{\pi}\,||\,(T^{-1})_{\sharp}\nu_{\pi})$) under the partial differential equation constraint (3) for the map $T$ (conversely, constraint (4) for the map $T^{-1}$).
@@
Triangular maps enable efficient numerical computation of a transport map and its inverse, which is also triangular. Moreover, triangular transport maps are well-suited for conditional density estimation and sampling [10, 24, 40, 2, 17]. Techniques for computational efficiency and scalability to high-dimensional problems include, among others, composition of low order maps [10] and the use of sparse triangular transport maps [29, 40, 2, 17]. For such sparse maps, any given map component $(T^{-1})_{i}$ in equation (17) may depend only on a proper subset (possibly empty) of the variables $\{y_{j}\}_{j=1}^{i-1}$ and on $y_{i}$. Dependence on $y_{i}$ is required. A diagonal map, in particular, has $x_{i}=(T^{-1})_{i}(y_{i})$ for each $i=1,\ldots,m$. The random variables that a given map component depends on is referred to as the set of active variables for the map component. Collectively, they define a sparsity pattern for the transport map. For a summary of these notions, see Figure 3.
@@
Figure 3: Pictorial representation of triangular transports maps. To simplify notation in Subfigures 3a3c, we denote $S\equiv T^{-1}$ for the transport map $T^{-1}$ from (17). In this example, $S:\mathbb{R}^{5}\rightarrow\mathbb{R}^{5}$. In each plot, the horizontal axis indexes the random variables $\{\mathcal{{Y}}_{j}\}_{j=1}^{5}$ and the vertical axis indexes the map components $\{S_{i}\}_{i=1}^{5}$. A square appears at the intersection of grid point $(j,i)$ if the $i$-th map component $S_{i}$ depends on the value $y_{j}$ of the $j$-th random variable $\mathcal{{Y}}_{j}$. For any given map component, the set of active variables is the set of random variables that the map component depends on. Collectively, the active variables define a sparsity pattern for the transport map.
Sparse triangular transport maps can be computed without a sparsity pattern specified a priori, using specially designed algorithms which have been shown to perform well for density estimation problems even when sample data is limited in amount [2]. Alternatively, a desired sparsity pattern could be specified explicitly and enforced when computing a transport map. As discussed in Section 2, one motivation for our research is density estimation from samples when available data is scarce. We hence consider sparse maps for our motivating application, as expanded upon in Section 5 further below. Next, we conclude the current Section 4 with some brief comments on density estimation via triangular transport maps and several illustrative examples.
@@ 4.2 Density Estimation from Data
Referring to equations (5)–(6), recall from Section 3 that we designate the probability measure $\nu_{\rho}$ as the reference measure and the probability measure $\nu_{\pi}$ as the target measure. The target density $\pi$ is unknown, but samples from the target measure $\nu_{\pi}$ are available. On the other hand, the reference measure $\nu_{\rho}$ is chosen to be known and so that it is amenable to computations and sampling. The reference measure is often selected to be the standard normal $\mathcal{N}(0,I)$, and that is our choice in all of our experiments.
@@
Given a set of samples from the target measure $\nu_{\pi}$, we use software libraries [28, 3] made available by the MIT Uncertainty Quantification (MUQ) Group [27] to compute a transport map (and its inverse) between the unknown target and reference measures. The transport map is computed as the solution of an optimization problem with objective function (18), using the Sample Average Approximation (SAA) algorithm. The reader is referred to [24] for full details.
@@
Upon computation of the transport map, the target density $\pi$ can be evaluated via equation (4). Given a (new) sample $y\sim\nu_{\pi}$, the value $\pi(y)$ can thus be obtained and used as needed, for instance for the classification problem from our motivating application, which was described in Section 2. The MUQ software libraries [28, 3] provide functionality for evaluating the target density, among many other capabilities.
@@
To describe an application of the measure transport framework, in Figure 4 are displayed computed approximations to density functions for univariate measures. The triangular transport maps for these approximations were computed using the TransportMaps library [28], version 2.0b3. The figure shows how the measure transport framework can be used for uni-modal and multi-modal data distributions alike. Further details are provided in the caption for Figure 4.
@@
Figure 4: Approximations to univariate density functions via triangular transport maps. In each of the Subfigures 4a and 4b, (i) the red curve is the reference density $\rho$ – a standard normal, (ii) a histogram of samples from the target density $\pi$ is depicted with black lines, (iii) the blue curve is the transport map (TM) approximation to the target density $\pi$, and (iv) the cyan curve is the density for the normal distribution $\mathcal{N}(\mu,\Sigma)$ with mean and standard deviation being that of the sample data. For 4a, the data set was created by drawing samples from a normal distribution $\mathcal{N}(\mu,\Sigma)$ with $\mu\neq 0$ and $\Sigma\neq 1$. The TM approximation to the target density agrees with that of the true density. The samples for the data set from 4b are from a multi-modal distribution. The resulting TM approximation to the target density reflects this and follows the data histogram more closely, as opposed to the density for $\mathcal{N}(\mu,\Sigma)$.
@@
Figure 5: Transport map (TM) approximations to class-conditional densities for the data set [20]. Subfigure 5a: Class 0 marginal distribution densities from data (far left) and three different TM approximations, with an increasing number of terms in the basis function expansions for the TM components. Clearly, the data density distribution is not Gaussian. The first TM approximation, with only one term in each basis function expansion, approximates the target density poorly. The accuracy of the approximations improves with an increased number of terms in the basis function expansions. Subfigure 5b: As in Subfigure 5a, but for class 1. Subfigure 5c: Left: reference probability density, which is a standard multivariate normal for all TM approximations. Afterwards: improved classification result as the accuracy of the density approximations increases. The top row confusion matrices are from training data and the bottom row ones are from inference data. Subfigure 5d: As the data is not Gaussian, a naive Bayes classifier (left) performs poorly. A support vector machine (middle) and neural network (right) classifier perform as well as the classification resulting from the most accurate TMs approximation (far right, Subfigures 5a-5c).
@@
Finally, transport map approximations to densities for a multivariate data set chosen from the UC Irvine Machine Learning Repository [18] appear in Figure 5. Although not from the biological domain, we selected the data set [20] because it yielded a good, non-contrived example for illustrating the benefits of a measure transport approach to density estimation. The data set was created from images taken from genuine and forged banknote-like specimens, creating two classes for classification problems. For the purpose of discussion, we identify them as class 0 and class 1 and compute approximations to the densities for each of the two classes. The triangular transport maps for the approximated densities were obtained using the AdaptiveTransportMaps library [3], which allows for computing adaptive parameterizations of transport maps [2].

### v4

@@ 3 Basics of Measure Transport
Recall that we are concerned with the problem of estimating an unknown probability density function characterizing some given data set of samples or observations, for instance, gene expression data. Denote the corresponding unknown probability measure by $\nu_{\psi}$ and call it the target measure. On the other hand, let $\nu_{\rho}$ be a probability measure amenable to computations, for example, the standard normal (i.e., Gaussian) measure $\mathcal{N}(0,I)$, so one can easily draw samples from it and evaluate its density $\rho$. Call this the reference measure. Now, let $\mathcal{{X}}=\left(\mathcal{{X}}_{1},\mathcal{{X}}_{2},\dots,\mathcal{{X}}_{m}\right)$ and $\mathcal{{Y}}=\left(\mathcal{{Y}}_{1},\mathcal{{Y}}_{2},\dots,\mathcal{{Y}}_{m}\right)$ be random vectors with probability distribution functions $F_{\mathcal{{X}}}$ and $F_{\mathcal{{Y}}}$, respectively, such that $\mathrm{d}{{F_{\mathcal{{X}}}}}=\nu_{\rho}$ and $\mathrm{d}{{F_{\mathcal{{Y}}}}}=\nu_{\psi}$. The random vectors $\mathcal{{X}}$, $\mathcal{{Y}}$ assume, respectively, values $x,y\in\mathbb{R}^{m}$ (also denoted by $x\sim\nu_{\rho}$ and $y\sim\nu_{\psi}$). Then if $X=\{x\in\mathbb{R}^{m}\,:\,x\sim\nu_{\rho}\}$ and $Y=\{y\in\mathbb{R}^{m}\,:\,y\sim\nu_{\psi}\}$ are sets of samples formed from the values of the random vectors $\mathcal{{X}}$ and $\mathcal{{Y}}$ (i.e., the elements in $X$ are distributed according to $\nu_{\rho}$, and those in $Y$ are distributed according to $\nu_{\psi}$) the transport map should satisfy $\nu_{\psi}(Y)\ =\ \nu_{\rho}(T^{-1}(Y))$ (by virtue of the measure transport equation (5)).
@@
Note that the set $Y$ consists of samples (usually taken from experiment) of the random vector $\mathcal{{Y}}$ distributed according to an a priori unknown target probability measure $\nu_{\psi}$, while the reference probability measure $\nu_{\rho}$ is a priori known (e.g., Gaussian) and the set $X$ is generated from samples of a random vector $\mathcal{{X}}$ distributed according to this known reference measure $\nu_{\rho}$. Thus, given the sets of samples $X$ and $Y$, our goal is to learn (i.e., compute) the transport map $T$ and use it later as a predictor for the values of $\mathcal{{Y}}$ and for evaluation of the target density $\psi$ (and therefore to serve as an estimator of the target measure $\nu_{\psi}$). Under the assumption that one can compute such a map $T$ transporting between the target and reference measures, the framework outlined herein and depicted in Figure 1 yields a method for approximating the desired unknown density function that characterizes the given sample data set $Y$. Then, the central problem is to compute, or learn, such a map $T$. This problem is further addressed in Section 4 below.
@@
Upon learning such a map $T$ from data, evaluation of the (unknown) density $\psi$ at a given (new) sample $y\sim\nu_{\psi}$ can be done via equation (7) as, by construction, all the functions on the right-hand side of equation (7) are known and can be evaluated. Moreover, because the reference measure is chosen to make sampling from it an easy and efficient task, sampling from the target measure can be done efficiently by drawing samples from the reference and transporting them to the target via the learned map $T$. Thus, as previously noted in Section 1, probabilistic modeling, inference, and statistical analysis in the (unknown) target space are enabled by such measure transport framework.
@@ 4 Transport Map Computation
To compute or, equivalently, learn the transport maps we seek, we use software [28, 2] made available by the MIT Uncertainty Quantification (MUQ) Group [27]. The reader is referred to the original publications [10, 24, 39, 3] for complete descriptions of the numerical methods and software libraries. A summary is provided next, to make the present manuscript self-contained.
Again, let $\nu_{\rho}$ and $\nu_{\psi}$ be two probability measures defined on the $\sigma$-algebra of Borel sets of $\mathbb{R}^{m}$, $m\geq 1$. Suppose that $T:\mathbb{R}^{m}\rightarrow\mathbb{R}^{m}$ is a transport map pushing forward $\nu_{\rho}$ to $\nu_{\psi}$. That is, it satisfies the equation (see the discussion surrounding equation (5) in Section 3)
$\nu_{\psi}$ $=$ $T_{\sharp}\nu_{\rho}$ (8)
for $T$. By definition,
$\nu_{\rho}$ $=$ $(T^{-1})_{\sharp}\nu_{\psi}$ (9)
@@
and one has that
$y_{i}$ $=$ $T_{i}(x_{1},x_{2},\ldots,x_{m}),\quad i=1,\ldots,m,$ (11)
where $y=(y_{1},y_{2},\ldots,y_{m})\sim\nu_{\psi}$ and $x=(x_{1},x_{2},\ldots,x_{m})\sim\nu_{\rho}$. In other words, $T$ maps any given sample $x$ from the reference measure $\nu_{\rho}$ to a sample $y=T(x)$ from the target measure $\nu_{\psi}$. Conversely, the inverse transport map $T^{-1}$ of $T$ has $m$ components
@@
and one has that
$x_{i}=(T^{-1})_{i}(y_{1},y_{2},\ldots,y_{m})\quad i=1,\ldots,m.$ (13)
(That is, $T^{-1}$ maps any given sample $y$ from the target measure $\nu_{\psi}$ to a sample $x=T^{-1}(y)$ from the reference measure $\nu_{\rho}$.)
A transport map $T$ is computed via the solution of a minimization problem with cost or objective function $C_{\rho}(T)$ generally defined as
$C_{\rho}(T)$ $=$ $\int_{\mathbb{R}^{m}}c(x,T(x))\,\nu_{\rho}(\mathrm{d}{{x}})\,,$ (14)
subject to the measure transport equation constraint (8) (or (9)). (More on this minimization problem in Section 4.1 below.) The structure of the map $T$ depends on $\rho$, $\psi$, and $c(\cdot,\cdot)$. Many types of maps (not necessarily optimal), or couplings, exist between measures [45, 36]. In this study we consider the use of triangular transport maps, as they enable practical numerical computation of a transport map and its inverse [24, 3].
@@ 4.1 Triangular Transport Maps
A transport map $T:\mathbb{R}^{m}\rightarrow\mathbb{R}^{m}$, satisfying the measure transport equation $\nu_{\psi}=T_{\sharp}\nu_{\rho}$, is triangular if
$y_{i}$ $=$ $T_{i}(x_{1},x_{2},\ldots,x_{i}),\quad i=1,\ldots,m$ (15)
(cf. equation (11)). In other words, the $i$-th component $T_{i}$ of a triangular transport map $T$ depends only on the values $\{x_{j}\}_{j=1}^{i}$ of the first $i$ random variables $\{\mathcal{{X}}_{j}\}_{j=1}^{i}$. Thus, the map components are
$T(x)$ $=$ $\left[\begin{array}[]{l}T_{1}(x_{1})\\ T_{2}(x_{1},x_{2})\\ T_{3}(x_{1},x_{2},x_{3})\\ \qquad\vdots\\ T_{m}(x_{1},x_{2},\ldots,x_{m})\end{array}\right]\,,$ (21)
from where one can observe the triangular structure of the map. Conversely, the inverse map $T^{-1}$ of $T$ is also triangular and, so,
$x_{i}$ $=$ $(T^{-1})_{i}(y_{1},y_{2},\ldots,y_{i}),\quad i=1,\ldots,m$ (22)
(cf. equation (13)).
To ensure (the required) positivity of the densities (6) and (7), each map component $T_{i}$ in equation (15) must be differentiable with respect to $x_{i}$, and it must hold that $\partial_{x_{i}}T_{i}>0$, for every $i=1,\ldots,m$. We refer to the class of maps with the latter property as the class of increasing triangular maps. Similarly, each component $(T^{-1})_{i}$ from equation (22) of the inverse map $T^{-1}$ of $T$ must satisfy said properties with respect to the variables $y_{i}$ [4, 24]. Furthermore, when both $\nu_{\rho}$ and $\nu_{\psi}$ (with densities, respectively, $\rho$ and $\psi$) are absolutely continuous with respect to the Lebesgue measure $\mathrm{d}{{x}}$ on $\mathbb{R}^{m}$, if in the cost function (14) we have $c(x,T(x))=||x-T(x)||^{2}$, where |||||$\cdot$||| | ⋅ | | is the Euclidean distance on $\mathbb{R}^{m}$, and $\nu_{\rho}$ is Gaussian, instead of the cost function $C_{\rho}(T)$ in (14) one may seek to minimize the relative entropy, or Kullback-Leibler divergence,
$\mathrm{D_{KL}}(\nu_{\psi}\,||\,\nu_{\rho})$ $=$ $\int_{\mathbb{R}^{m}}\psi\log(\psi/\rho)\,\mathrm{d}{{x}}\ =\ \mathbb{E}_{\psi}(\log(\psi/\rho))\,,$ (23)
as $C_{\rho}(T)\leq 2A\,\mathrm{D_{KL}}(T_{\sharp}\nu_{\rho}\,||\,\nu_{\rho})$, where $A>0$ is a constant [31, 19, 4]. Under the assumption that the map $T$ is differentiable or even smooth, one may thus look at the problem of learning a transport map $T$ as a solution of the optimization problem
@@
(conversely, $\arg\min_{T}\mathrm{D_{KL}}(\nu_{\psi}\,||\,(T^{-1})_{\sharp}\nu_{\psi})$) subject to the partial differential equation constraint (6) for the map $T$ (conversely, constraint (7) for the map $T^{-1}$).
@@
Triangular maps enable efficient numerical computation of a transport map and its inverse, which is also triangular. Moreover, triangular transport maps are well-suited for conditional density estimation and sampling [10, 24, 39, 3, 17]. Techniques for computational efficiency and scalability to high-dimensional problems include, among others, composition of low order maps [10] and the use of sparse triangular transport maps [29, 39, 3, 17]. For such sparse maps, any given map component $(T^{-1})_{i}$ in equation (22) may depend only on a proper subset (possibly empty) of the variables $\{y_{j}\}_{j=1}^{i-1}$ and on $y_{i}$. Dependence on $y_{i}$ is required. A diagonal map, in particular, has $x_{i}=(T^{-1})_{i}(y_{i})$ for each $i=1,\ldots,m$. The random variables that a given map component depends on is referred to as the set of active variables for the map component. Collectively, the active variables define a sparsity pattern for the transport map. For a summary of these notions, see Figure 3. We note that the notion of sparsity pattern characterizes the triangular transport map and reflects the dependence structure of the target probability measure (i.e., by virtue of being triangular, it represents dependence relations between the random variables jointly distributed according to the target probability measure). Thus, “sparsity” here does not refer to the number of samples available for learning a triangular transport map.
@@
Figure 3: Pictorial representation of triangular transport maps sparsity patterns. To simplify notation in Subfigures 3a3c, we denote $S\equiv T^{-1}$ for the transport map $T^{-1}$ from (22). In this example, $S\!:\!\mathbb{R}^{5}\!\rightarrow\!\mathbb{R}^{5}$. In each plot, the horizontal axis indexes the random variables $\{\mathcal{{Y}}_{j}\}_{j=1}^{5}$ and the vertical axis indexes the map components $\{S_{i}\}_{i=1}^{5}$. A square appears at the intersection of grid point $(j,i)$ if the $i$-th map component $S_{i}$ depends on the value $y_{j}$ of the $j$-th random variable $\mathcal{{Y}}_{j}$. For any given map component, the set of active variables is the set of random variables that the map component depends on. Collectively, the active variables define a sparsity pattern for the transport map.
As explained in detail in the original publications [24, 3] and summarized in Appendix A of the present paper, the algorithms we use and build upon rely on finite-dimensional approximations to the functional representations for the transport map components, which may be described generally to depend on functions $f_{i}$ of the form
[…]
where, for each $i=1,\ldots,m$,
  • $\Lambda^{(i)}\subset\mathbb{Z}_{+}^{i}$ is a finite set of multi-indices, which identifies the active variables of the $i$-th transport map component and therefore encodes information about the sparsity pattern of the transport map,
  • $c_{\alpha^{(i)}}\in\mathbb{R}$ are unknown real-valued coefficients, which are to be learned from the data, and
  • ~$\widetilde{\mathcal{H}}_{\alpha^{(i)}}:\mathbb{R}^{i}\rightarrow\mathbb{R}$ are suitably defined basis functions (refer to Appendix A and [24, 3]).
Sparse triangular transport maps can be learned by explicitly specifying a sparsity pattern for the transport map and enforcing the pattern (via the sets $\Lambda^{(i)}$ of multi-indices from equations (25)) when learning the transport map. Alternatively, (sparse) triangular transport maps may be learned without a sparsity pattern being specified a priori, by learning the sets $\Lambda^{(i)}$ of multi-indices in equations (25) along with the unknown coefficients $c_{\alpha^{(i)}}$, using specifically designed algorithms which have been shown to perform well for density estimation problems that considered a wide range of values (mainly $O(100)$ or larger) for the number of training samples used to learn the transport maps [3].
Recall from Sections 1.1 and 2 that one motivation for our research is density estimation from data when the number of samples available for learning the model (within this context, learning a transport map) is limited in amount (more precisely, of order $O(1)$$O(10)$ number of samples). For the purpose we consider the use of triangular transport maps, first by explicitly enforcing transport map sparsity patterns (Section 5.1) and second by not enforcing transport map sparsity patterns a priori at all (Section 5.2). In order to avoid possible confusion, we emphasize again that the sparsity pattern of a triangular transport map is a property of the target probability measure and not a property related to the number of samples used for transport map learning. Next, we conclude the current Section 4 with some brief comments on density estimation via triangular transport maps and several illustrative examples.
@@ 4.2 Density Estimation from Data
Referring to equations (8)–(9), recall from Section 3 that we designate the probability measure $\nu_{\rho}$ as the reference measure and the probability measure $\nu_{\psi}$ as the target measure. The target density $\psi$ of interest to us is unknown, but samples from the target measure $\nu_{\psi}$ are available. On the other hand, the reference measure $\nu_{\rho}$ is known and purposefully chosen so that its density $\rho$ can be evaluated in an efficient manner and so that sampling from $\nu_{\rho}$ can be done efficiently as well. The reference measure $\nu_{\rho}$ is often selected to be the standard normal (i.e., Gaussian) measure $\mathcal{N}(0,I)$, and that is our choice in all of our computational experiments.
@@
Given a set of samples from the unknown target measure $\nu_{\psi}$, we use software libraries [28, 2] made available by the MIT Uncertainty Quantification (MUQ) Group [27] to learn a transport map (and its inverse) between the target and reference measures. The transport map is learned as the solution of an optimization problem with objective function minimizing (23), using the Sample Average Approximation (SAA) algorithm. It is outside the scope of this paper to cover the details pertaining to the solution of the optimization problem (although we provide a summary of some aspects in Appendix A), but the interested reader is referred to [24, 3] for a full account.
@@
Recall from Section 3 that upon learning a transport map $T$ that “couples” the reference and target measures, the target density $\psi$ can be evaluated via equation (7) (which, within the context of this paper, corresponds to equation (3)) and, so, given a (new) sample $y$, the value $\psi(y)$ can thus be obtained and used as needed. Hence our density estimation problem is solved by the measure transport framework. The MUQ software libraries [28, 2] used in our computations provide functionality for evaluating the target density $\psi$ and sampling from the target measure $\nu_{\psi}$, among many other capabilities.
@@
To describe an application of the measure transport framework, in Figure 4 are displayed transport map computed approximations to density functions for univariate measures. The triangular transport maps for these approximations were learned using the TransportMaps library [28], version 2.0b3. The figure shows how the measure transport framework can be used for uni-modal and multi-modal data distributions alike. In each case, we learned the transport maps using synthetic data sets (i.e., data sets generated by us). In Subfigure 4a, the target is a Gaussian measure $\mathcal{N}(\mu,\Sigma)$ with mean $\mu\neq 0$ and standard deviation $\Sigma\neq 1$, and so we know the formula $\psi_{true}(y)=(\Sigma)^{-1}(2\pi)^{-\frac{1}{2}}\exp(-\frac{1}{2}(\frac{y-\mu}{\Sigma})^{2})$ for the true target density. We can thus verify that the transport map approximation (7) to the target $\psi$ (blue curve in Subfigure 4a) agrees with that of the true density $\psi_{true}(y)$ (cyan curve in Subfigure 4a). On the other hand, the target density $\psi$ in Subfigure 4b is multi-modal and, hence, the density for a Gaussian measure (cyan curve in Subfigure 4b) with mean and standard deviation being that of the sample data is a poor approximation to the data histogram (black curve in Subfigure 4b). The transport map approximation to the target density $\psi$ (blue curve in Subfigure 4b) captures the multi-modality and better approximates the data histogram (black curve in Subfigure 4b). Further details are provided in the caption for Figure 4.
@@
Figure 4: Approximations to univariate density functions via triangular transport maps. In each of the Subfigures 4a and 4b, (i) the red curve is the reference density $\rho$ – which is the density of a standard normal (i.e., Gaussian) measure $\mathcal{N}(0,1)$, (ii) a histogram of samples from the target measure $\nu_{\psi}$ is depicted with black lines, (iii) the blue curve is the transport map (TM) approximation to the target density $\psi$, and (iv) the cyan curve is the density for the normal (i.e., Gaussian) measure $\mathcal{N}(\mu,\Sigma)$ with mean and standard deviation being that of the sample data. For Subfigure 4a, the data set was created by drawing samples from a normal (i.e., Gaussian) measure $\mathcal{N}(\mu,\Sigma)$ with $\mu\neq 0$ and $\Sigma\neq 1$. As can be seen, the TM approximation to the target density (blue curve) agrees with that of the true density (cyan curve). The samples for the data set from Subfigure 4b were drawn from a multi-modal target measure $\nu_{\psi}$. The resulting TM approximation to the target density (blue curve) reflects this and follows the data histogram (black curve) more closely, as opposed to the density for $\mathcal{N}(\mu,\Sigma)$ (cyan curve).
@@
Figure 5: Transport map (TM) approximations to class-conditional densities for the data set [20]. Subfigure 5a: Class 0 marginal distribution densities from data (1st subplot, far left) and three different TM approximations (2nd through 4th subplots), with an increasing number of terms in the basis function expansions (47) for the TM components. Clearly, the data density distribution (1st subplot in Figure 5a) is not Gaussian. The first TM approximation (2nd subplot in Figure 5a), with only one term in each basis function expansion (47), approximates the target density (1st subplot in Figure 5a) poorly. The accuracy of the approximations improves (3rd and 4th subplots in Figure 5a, compared to 1st subplot in Figure 5a) with an increased number of terms in the basis function expansions (47). Subfigure 5b: As in Subfigure 5a, but for class 1. Subfigure 5c: Left: reference probability density, which is a standard multivariate normal for all TM approximations. Afterwards: confusion matrices showing improved classification result as the accuracy of the TM density approximations increases. The top row confusion matrices are from training data and the bottom row ones are from inference data. Subfigure 5d: As the data is not Gaussian, a naive Bayes classifier (left) performs poorly. A support vector machine (middle) and neural network (right) classifier perform as well as the classification resulting from the most accurate TMs approximation (far right, Subfigures 5a-5c).
@@
Finally, transport map approximations to densities for a multivariate data set chosen from the UC Irvine Machine Learning Repository [18] appear in Figure 5. Although not from the biological domain, we selected the data set [20] because it yielded a good, non-contrived example for illustrating the benefits of a measure transport approach to density estimation. The data set [20] was created by its author from images taken from genuine and forged banknote-like specimens, creating two classes for classification problems. For the purpose of discussion, we identify the classes as class 0 and class 1 and compute approximations to the densities for each of the two classes. The triangular transport maps for the approximated densities were obtained using the AdaptiveTransportMaps library [2], which allows for computing adaptive parameterizations of transport maps [3].